OpenStudy (anonymous):

Two cars leave an intersection at the same time.One heads due south and other heads due east. Later, the two cars are exactly 50 mi apart. The car headed south has traveled 10 mi farther the car going east. How far has the car going east traveled?

OpenStudy (anonymous):

Okay. Let x be the distance that the car moving East has traveled. Then x+10 is the distance that the car moving South has traveled. Solve the following for x: \[x^2 + (x + 10)^2 = 50^2\] \[x = 30\]

OpenStudy (anonymous):

Hope this helped! Have a great day! @Yishai1

OpenStudy (anonymous):

Thank you